OpenStudy (anonymous):

Which of the following equations would generate the sequence 4,7,10,13... A. an=4n+3 B. an=4+3(n-1) C. an=3n+4 D. an=4+7(n-1)

OpenStudy (theeric):

Ah! It probably looks like \(a_n=4n+3\) and so on, right? Well first, do you see a pattern in the numbers?

OpenStudy (atlas):

One way is to put values of n =0,1,2 ,3............and check which equation gives the correct sequence

OpenStudy (atlas):

no eric you are wrong

OpenStudy (theeric):

@atlas has a good idea, too!

OpenStudy (theeric):

I am wrong?

OpenStudy (anonymous):

well i know it increases by 3 so i thought it was A also but when i did 4(2)+3 i got 11 instead of 10

OpenStudy (anonymous):

so thats where im stuck

OpenStudy (theeric):

Then is \(a_n\) really \(a\times n\)? I was NOT supposing I gave the answer. I was questioning the formatting, and started with the first option and said, "and so on."

OpenStudy (atlas):

Another way is to see that each number in the sequence increases by 3 and the first number is 4. so sequence is : 4, 4+3, 4+3+3, 4+3+3 and so on

OpenStudy (theeric):

Please, @atlas , don't immediately assume someone is wrong without questioning them or what you think is right - it leads to a less educational environment, in my opinion... But It was probably just a miscommunication! :)

OpenStudy (atlas):

So you see it should be 4+multiple of 3 i.e 4+3n

OpenStudy (atlas):

I am sorry theEric: I hope you see your mistake

OpenStudy (anonymous):

it isnt a x n it is a sub n

OpenStudy (theeric):

So now we have two ways to look at it! 1. Use the options and see which one works. 2. Try to look at the pattern and form an expression for \(a_n\). Both are just as good!

OpenStudy (anonymous):

thank you both! :)

OpenStudy (anonymous):

so 4+3n would create the same numbers in the arithmetic sequence?

OpenStudy (atlas):

Of course...........why don't you try out and see

OpenStudy (gorv):

an=4+3(n-1)

OpenStudy (gorv):

this is the ans

ganeshie8 (ganeshie8):

4 7 10 13 3 3 3 \(a_n = 4+3(n-1)\)

OpenStudy (gorv):

put n=1,2,3,4,5,6,7.........

OpenStudy (atlas):

@gorv: yeah it depends on whether you start from n=0 or n=1

OpenStudy (atlas):

I am a digital guy so I start counting from 0 :P

ganeshie8 (ganeshie8):

lol mee too a binary guy :) n starts from 1, unless otherwise specified

OpenStudy (gorv):

yeah buddyyy

OpenStudy (atlas):

yeah if you count from n=1 -> the answer is 1+3n as others have pointed out :P

OpenStudy (anonymous):

sorry i was eating lunch but thank you so much guys! :)

OpenStudy (anonymous):

so the answer would be B. an=4+3(n-1)? :o

OpenStudy (anonymous):

@atlas

OpenStudy (atlas):

yeah

OpenStudy (anonymous):

thank you, just wanted to confirm :)

OpenStudy (atlas):

@ganeshie8 is right

OpenStudy (theeric):

I agree! :)

OpenStudy (anonymous):

woohoo! @ganeshie8 is awesome & so are you guys! I wish i could give all of you guys medals :( lol

OpenStudy (theeric):

Haha, glad to help!

OpenStudy (anonymous):

so for next time you dont start at 0?

ganeshie8 (ganeshie8):

i can see theEric was trying to help u by giving u a wrong example, and then proving to u why it wudnt work... and then we entered and spoiled the thing :)

OpenStudy (anonymous):

hehe all of you guys were a great help! i really appreciate it :)

OpenStudy (atlas):

I still count from 0 It is a habit . :P

OpenStudy (anonymous):

Eric helped me see why my first choice was wrong, because i had intially thought A was right :)

ganeshie8 (ganeshie8):

aww thnks, never start at 0, unless its specified explicitly in the question. sequences start from 1. in general humans count from 1, and machines count from 0 ! :P

OpenStudy (theeric):

Haha, I gave no example, sorry! Miscommunication! "Ah! It probably looks like an=4n+3 and so on, right? Well first, do you see a pattern in the numbers?" That was as compared to "A. an=4n+3". It is not \(an=4n+3\). It is \(a_n=4n+3\). The an=4n+3 was just the lack of formatting! But you can test to eliminate options like that :) Thanks!

OpenStudy (anonymous):

makes sense! thanks @ganeshie8 :)

OpenStudy (theeric):

:) Take care all! Sorry for the confusion that I introduced!

OpenStudy (atlas):

All arithmetic expressions follow the same pattern : first term + difference * (n-1) 4 + 3 * (n-1)

OpenStudy (atlas):

Check with this : 1,6,11,16.21.......... ??

ganeshie8 (ganeshie8):

ahh start wid n=1 4, 7, 10, 13, 16,....

OpenStudy (theeric):

I see you guys are still talking about this, so I'll share a link with you. It might be a nice part of your discussion, I don't know! http://www.algebralab.org/lessons/lesson.aspx?file=Algebra_ArithSeq.xml